Express the function f: A—R. f(x) = x2 – 1. where A = { -4, 0, 1, 4) as a set of ordered pairs.

Grade:12th pass

1 Answers

SJ
3 years ago
Welcome to Askiitians

Given,

A = {-4, 0, 1, 4}

f(x) = x2– 1

f(-4) = (-4)2– 1 = 16 – 1=15

f(0) = (0)2– 1 = -1

f(1) = (1)2– 1 = 0

f(4) = (4)2– 1 = 16 – 1 =15

Therefore, the set of ordered pairs = {(-4, 15), (0, -1), (1, 0), (4, 15)}

Thanks